# Multifractality and Hyperuniformity in Quasicrystalline Bose-Hubbard Models with and without Disorder

### v2

@@ Abstract
Clarifying similarities and differences in physical properties between crystalline and quasicrystalline systems is one of central issues in studying quasicrystals. To contribute to this, we apply multifractal and hyperuniform analyses to nonuniform spatial patterns in the Bose-Hubbard model on the Penrose and Ammann-Beenker tilings. Based on the mean-field approximation, we obtain real-space distributions of local superfluid amplitude and bosonic condensate. In both Mott insulating and superfluid phases, the distributions are hyperuniform. Analyzing the order metric that quantifies the complexity of nonuniform spatial patterns, we find that both quasicrystals show a significant increase of the order metric at a phase boundary between the Mott insulating and superfluid phases, in stark contrast to the case of a periodic square lattice. Our results suggest that hyperuniformity is a useful concept to differentiate between crystalline and quasicrystalline bosonic systems. The order metric clarifies if the distribution of a physical quantity reflects the point distribution or not, and quantifies how complex the distribution is in comparison with the point distribution. Moreover, we introduce on-site random potentials into these quasicrystalline Bose-Hubbard models, leading to a Bose glass phase. Contrary to the Mott insulating and superfluid phases, we find that the Bose glass phase is multifractal. The same multifractality appears on a Bose glass phase in the periodic square lattice. Therefore, multifractality is common in a Bose glass phase irrespective of the periodicity of systems.
@@ 1 Introduction
In contrast to states in fermion systems, multifractality and hyperuniformity in boson systems have not been elucidated, although bosonic quasicrystalline systems have been examined experimentally [20, 21, 22, 23] and theoretically [24, 25, 26] by using ultra-cold atoms in optical quasicrystals. Such boson systems can be described by a Bose-Hubbard model, which exhibits Mott insulating phase (superfluid phase) when on-site Coulomb interaction (hopping interaction) is dominant [24]. It has not been clarified whether spatial distribution of physical quantities in those phases exhibits either hyperuniform or multifractal.
@@ 2.2 Perpendicular space
We consider the Bose-Hubbard model (2) on a Penrose tiling that is a covering of the plane by non-overlapping rhombi as shown in Fig. 1(a). Since a Penrose tiling forms a quasicrystalline structure, we call structures formed by a Penrose titling Penrose quasicrystals in this paper. The vertices of these quasicrystals are obtained by projection from a five-dimensional hypercubic lattice onto a two-dimensional plane. For the direction perpendicular to this projection, we consider another projection from a five-dimensional hypercubic lattice onto a three-dimensional plane, which results in the perpendicular-space representation as shown in Fig. 1(b). Any vertex is located at one of the four pentagonal planes at $\tilde{z}=0,$ $1/\sqrt{2},$ 2/$2/\sqrt{2},$ and $3/\sqrt{2}$. We show two of the four planes for $\tilde{z}=1/\sqrt{2}$ (left panel) and 0 (right panel) in Fig. 1(c), where the colors distinguish a coordination number $Z_{i}$ defined as the number of vertices that are connected with the $i$th vertex along links. The black lines separate vertices with different $Z_{i}$, resulting in the formation of star shapes in the pentagonal planes. The average of $Z_{i}$ for the entire system is $\overline{Z_{i}}=4$ in Penrose quasicrystals, which agrees with that of square lattice. Note that the pentagonal planes at $\tilde{z}=3/\sqrt{2}$ and $\tilde{z}=2/\sqrt{2}$ are identical to the planes at $\tilde{z}=0$ and $\tilde{z}=1/\sqrt{2}$, respectively, with reversing them upside down. This allows us to focus on only two planes at $\tilde{z}=0$ and $1/\sqrt{2}$.
@@ 2.3 Multifractality
The definition of multifractality depends on a spatial dimension. In this paper, we concentrate on the multifractality in two dimensions. We introduce a circular window of radius $R$ and consider a $R$-dependent quantity given by
@@ 3.2 Near phase boundary
We focus on a phase boundary between the Mott insulating and superfluid phases. At the vicinity of the phase boundary, the system may show an exotic phase such as Bose glass [37, 38, 26]. The system is in the Bose glass phase when superfluid vertices are randomly distributed. On the other hand, the system is in the superfluid phase when the fraction of superfluid vertices is large. We note that the Bose glass phase has been observed in a system of ultra-cold atoms in optical quasicrystals [21, 23].
@@ 3.2.3 Comparison between crystals and quasicrystals
Figure 7(d) shows a relationship between $a$ and $Z_{i}$ in a Penrose quasicrystal. The blue circles represent mean values of $a$ with respect to vertices with the same $Z_{i}$. As these mean values are almost on a straight line, we show a linear fitting of them by a blue-dashed line. It implies that $|\braket{\hat{b}_{i}}|$ near the phase transition point is almost proportional to the number of linked vertices $Z_{i}$. This is reasonable because the mean field should be larger for a vertex with larger $Z_{i}$ [see Eq. (2)]. A value of $a$ for a square lattice is indicated by a green-dashed line, which intersects with the blue-dashed line between $Z_{i}=4$ and 5. Although the average value of $Z_{i}$ in Penrose quasicrystals is 4 which is the same as in square lattice, we find that $a$ for a square lattice is larger than $a$ for $Z_{i}=4$ in a Penrose quasicrystal. We show a result for Ammann-Beenker quasicrystal in Fig. 7(e), where we also find a linear dependence of $Z_{i}$ on $a$. The intersection of the blue-dashed and green-dashed lines is between $Z_{i}=4$ and 5 in an Ammann-Beenkwe quasicrystal.
@@ 5 Conclusion
In summary, we have investigated multifractal or hyperuniform properties in the Bose-Hubbard model on the Penrose and Ammann-Beenker tilings with and without disorder. We have found that, without disorder, the system is hyperuniform in both Mott insulating and superfluid phases. From the enhancement of order metric, we have found that the complexity of the system increases near the phase boundary of the two phases. This is in contrast to a constant order metric in square lattice. A wider variety of vertices in quasicrystals allows more intricate distribution of a physical quantity, which results in the larger order metric than in crystals. The difference between quasicrystals and crystals is more significant as approaching a phase boundary because of an increase in a typical correlation length of a physical quantity. Examining a critical hopping strength for generating superfluid vertex near the phase boundary, we have found that the distribution of a physical quantity near a phase boundary reflects the quasiperiodicity of the point distribution in quasicrystals. Examining a critical hopping strength for generating superfluid vertex near the phase boundary, we have attributed the origin of enhanced order metric to quasiperiodicity in the systems. This contrasting behavior between crystalline and quasicrystalline systems can be confirmed by using ultra-cold atoms if one measures the distribution of local superfluid amplitude in optical lattice [20, 21, 22, 23].
@@ Acknowledgements.
This work was supported by JST SPRING (Grant No. JPMJSP2151) and the Japan Society for the Promotion of Science, KAKENHI (Grant No. JP19H05821).

### v3

@@ Abstract
Clarifying similarities and differences in physical properties between crystalline and quasicrystalline systems is one of central issues in studying quasicrystals. To contribute to this, we apply multifractal and hyperuniform analyses to nonuniform spatial patterns in the Bose-Hubbard model on the Penrose and Ammann-Beenker tilings. Based on the mean-field approximation, we obtain real-space distributions of local superfluid amplitude and boson density. In both Mott insulating and superfluid phases, the distributions are hyperuniform. Analyzing the order metric that quantifies the complexity of nonuniform spatial patterns, we find that both quasicrystals show a significant increase of the order metric at a phase boundary between the Mott insulating and superfluid phases, in stark contrast to the case of a periodic square lattice. Our results suggest that hyperuniformity is a useful concept to differentiate between crystalline and quasicrystalline bosonic systems. The order metric clarifies if the distribution of a physical quantity reflects the point distribution or not, and quantifies how complex the distribution is in comparison with the point distribution. Moreover, we introduce on-site random potentials into these quasicrystalline Bose-Hubbard models, leading to a Bose glass phase. Contrary to the Mott insulating and superfluid phases, we find that the Bose glass phase is multifractal. The same multifractality appears on a Bose glass phase in the periodic square lattice. Therefore, multifractality is common in a Bose glass phase irrespective of the periodicity of systems.
@@ 1 Introduction
In contrast to states in fermion systems, multifractality and hyperuniformity in boson systems have not been elucidated, although bosonic quasicrystalline systems have been examined experimentally [20, 21, 22, 23] and theoretically [24, 25, 26] by using ultracold atoms in optical quasicrystals. Such boson systems can be described by a Bose-Hubbard model, which exhibits Mott insulating phase (superfluid phase) when on-site Coulomb interaction (hopping interaction) is dominant [24]. It has not been clarified whether spatial distribution of physical quantities in those phases exhibits either hyperuniform or multifractal.
@@ 2.2 Perpendicular space
We consider the Bose-Hubbard model (2) on a Penrose tiling that is a covering of the plane by non-overlapping rhombi as shown in Fig. 1(a). Since a Penrose tiling forms a quasicrystalline structure, we call structures formed by a Penrose titling Penrose quasicrystals in this paper. The vertices of these quasicrystals are obtained by projection from a five-dimensional hypercubic lattice onto a two-dimensional plane. For the direction perpendicular to this projection, we consider another projection from a five-dimensional hypercubic lattice onto a three-dimensional space, which results in the perpendicular-space representation as shown in Fig. 1(b). Any vertex is located at one of the four pentagonal planes at $\tilde{z}=0,$ $1/\sqrt{2},$ 2/$2/\sqrt{2},$ and $3/\sqrt{2}$. We show two of the four planes for $\tilde{z}=1/\sqrt{2}$ (left panel) and 0 (right panel) in Fig. 1(c), where the colors distinguish a coordination number $Z_{i}$ defined as the number of vertices that are connected with the $i$th vertex along links. The black lines separate vertices with different $Z_{i}$, resulting in the formation of star shapes in the pentagonal planes. The average of $Z_{i}$ for the entire system is $\overline{Z_{i}}=4$ in Penrose quasicrystals, which agrees with that of square lattice. Note that the pentagonal planes at $\tilde{z}=3/\sqrt{2}$ and $\tilde{z}=2/\sqrt{2}$ are identical to the planes at $\tilde{z}=0$ and $\tilde{z}=1/\sqrt{2}$, respectively, with reversing them upside down. This allows us to focus on only two planes at $\tilde{z}=0$ and $1/\sqrt{2}$.
@@ 2.3 Multifractality
The definition of multifractality depends on a spatial dimension. In this paper, we concentrate on the multifractality in two dimensions. We introduce a circular window of radius $R$ and consider an $R$-dependent quantity given by
@@ 3.2 Near phase boundary
We focus on a phase boundary between the Mott insulating and superfluid phases. At the vicinity of the phase boundary, the system may show an exotic phase such as Bose glass [37, 38, 26]. The system is in the Bose glass phase when superfluid vertices are randomly distributed. On the other hand, the system is in the superfluid phase when the fraction of superfluid vertices is large. We note that the Bose glass phase has been observed in a system of ultracold atoms in optical quasicrystals [21, 23].
@@ 3.2.3 Comparison between crystals and quasicrystals
Figure 7(d) shows a relationship between $a$ and $Z_{i}$ in a Penrose quasicrystal. The blue circles represent mean values of $a$ with respect to vertices with the same $Z_{i}$. As these mean values are almost on a straight line, we show a linear fitting of them by a blue-dashed line. It implies that $|\braket{\hat{b}_{i}}|$ near the phase transition point is almost proportional to the number of linked vertices $Z_{i}$. This is reasonable because the mean field should be larger for a vertex with larger $Z_{i}$ [see Eq. (2)]. A value of $a$ for a square lattice is indicated by a green-dashed line, which intersects with the blue-dashed line between $Z_{i}=4$ and 5. Although the average value of $Z_{i}$ in Penrose quasicrystals is 4 which is the same as in square lattice, we find that $a$ for a square lattice is larger than $a$ for $Z_{i}=4$ in a Penrose quasicrystal. We show a result for Ammann-Beenker quasicrystal in Fig. 7(e), where we also find a linear dependence of $Z_{i}$ on $a$. The intersection of the blue-dashed and green-dashed lines is between $Z_{i}=4$ and 5 in an Ammann-Beenker quasicrystal.
@@ 5 Conclusion
In summary, we have investigated multifractal or hyperuniform properties in the Bose-Hubbard model on the Penrose and Ammann-Beenker tilings with and without disorder. We have found that, without disorder, the system is hyperuniform in both Mott insulating and superfluid phases. From the enhancement of order metric, we have found that the complexity of the system increases near the phase boundary of the two phases. This is in contrast to a constant order metric in square lattice. A wider variety of vertices in quasicrystals allows more intricate distribution of a physical quantity, which results in the larger order metric than in crystals. The difference between quasicrystals and crystals is more significant as approaching a phase boundary because of an increase in a typical correlation length of a physical quantity. Examining a critical hopping strength for generating superfluid vertex near the phase boundary, we have found that the distribution of a physical quantity near a phase boundary reflects the quasiperiodicity of the point distribution in quasicrystals. Examining a critical hopping strength for generating superfluid vertex near the phase boundary, we have attributed the origin of enhanced order metric to quasiperiodicity in the systems. This contrasting behavior between crystalline and quasicrystalline systems can be confirmed by using ultracold atoms if one measures the distribution of local superfluid amplitude in optical lattice [20, 21, 22, 23].
@@ Acknowledgements.
This work was supported by the Natural Sciences and Engineering Research Council of Canada, JST SPRING (Grant No. JPMJSP2151) and the Japan Society for the Promotion of Science, KAKENHI (Grant No. JP19H05821).